OpenStudy (shesolitt):

I WILL MEDAL AND AND !! PLS HELP Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form. y = x + 8 y = x − 8 y = −5x − 10 y = −5x + 20

OpenStudy (shesolitt):

3 what? @AceSpeedFighter you mean C?

OpenStudy (shesolitt):

can you help me with one more pls? @AceSpeedFighter

OpenStudy (shesolitt):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form. y + 2 = 4(x – 3) y – 3 = 4(x + 2) x – 3 = 4(y + 2) x + 2 = 4(y – 3)

pooja195 (pooja195):

Hi there! :) Let's do one question at a time. For the original question (#1) you would start off by finding the slope. Use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \]

pooja195 (pooja195):

Go ahead and find the slope.

OpenStudy (shesolitt):

I never deleted them. I actually thought u did. idk whats happening @AceSpeedFighter

OpenStudy (acespeedfighter):

:/